OpenStudy (anonymous):

Assume f(x)=-2x+8 and g(x)=3x. What is the value of (g o f)(3)?

OpenStudy (anonymous):

gof(x)=g(f(x))=g(-2x+8)=3(8-2x)=24-6x gof(3)=24-6*3=24-18=6